# Closed-Form Training Dynamics Reveal Learned Features and Linear Structure in Word2Vec-like Models

### v2

@@ 4 Linear Structure in Latent Space
Figure 2 reveals that, for QWEMs, the “natural” basis of the learning dynamics is simply the eigenbasis of ${\bm{M}}^{\star}$. Figure 3 suggests that this basis already encodes concepts interpretable to humans. These are the fundamental features learned by the model: each word embedding is naturally expressed as a linear combination of these orthogonal latents. Both early stopping and small embedding dimension regularize the model by constraining the number (but not the character) of available latents.
We may conclude that natural language contains linear semantic structure in its co-occurrence statistics, and that it is easily extracted by word embedding algorithms. It is not unreasonable to expect, then, that other semantic concepts may be encoded as linear subspaces. This idea is the linear representation hypothesis, and it motivates modern research in more sophisticated language models, including representation learning (Park et al., 2023; Jiang et al., 2024; Wang et al., 2024), mechanistic interpretability (Li et al., 2023; Nanda et al., 2023; Lee et al., 2024), and LLM alignment (Lauscher et al., 2020; Zou et al., 2023; Li et al., 2024). To make these efforts more precise, it is important to develop a quantitative understanding of these linear representations in simple models. Our 1 provides a new lens for this analysis: we may gain insight by studying the properties of ${\bm{M}}^{\star}$.
In word embedding models, a natural category of abstract linear representations are the difference vectors between semantic pairs, e.g., $\{{\bm{r}}^{(n)}\}_{n}=\{\mathbf{man}-\mathbf{woman},\mathbf{king}-\mathbf{queen},\dots\}$ for gender binaries. If the ${\bm{r}}^{(n)}$ are all approximately equal, then the model can effectively complete analogies via vector addition (see Section A.4). Following Ilharco et al. (2022), we call the ${\bm{r}}^{(n)}$ task vectors.
@@
Task vectors derived from the analogy dataset are neither strongly aligned with a single model eigen-feature, nor are they random vectors. Instead, they exhibit localization: a handful of the top eigen-features are primarily responsible for the task vector. In some cases, these dominant eigen-features are interpretable and clearly correlate with the abstract semantic concept associated with the task vector (Figure 4).
@@
To study the geometry of the task vectors within a class of semantic binaries, we consider stacking task vectors to produce the matrix ${\bm{R}}_{d}\in\mathbb{R}^{N\times d}$, where $N$ is the number of word pairs and $d$ is the embedding dimension. We note that ${\bm{R}}_{d}$ can be computed in closed form using our 1. If all the task vectors align (as desired for analogy completion), then ${\bm{R}}_{d}$ will be a rank-1 matrix; if the task vectors are all unrelated, then ${\bm{R}}_{d}$ will have a broad spectrum. This observation suggests that a useful object to study is the empirical spectrum of the Gram matrix ${\bm{G}}_{d}\coloneqq{\bm{R}}_{d}{\bm{R}}_{d}^{\top}\in\mathbb{R}^{N\times N}$.
We find that this spectrum is very well-described by the spiked covariance model, a well-known distribution of random matrices. In the model, one studies the spectrum of ${\bm{Z}}=\bm{\Xi}{\bm{\Xi}}^{\top}+\mu{\bm{a}}{{\bm{a}}}^{\top}$, where $\bm{\Xi}$ is a random matrix with i.i.d. mean-zero entries with variance $\sigma^{2}$ and $\mu{\bm{a}}{{\bm{a}}}^{\top}$ is a deterministic rank-1 perturbation with strength $\mu$. If $\mu$ is sufficiently large compared to $\sigma$, then in the asymptotic regime the spectrum of ${\bm{Z}}$ is known to approach the Marchenko-Pastur distribution with a single outlier eigenvalue (Baik et al., 2005). We consistently observe this structure in the real empirical data, across both model dimension and semantic families (Figure 4). We note that $N\approx 30$ is fairly small, and so it is somewhat surprising that results from the asymptotic regime visually appear to hold.
@@
Figure 4: Models build linear representations from a few informative and many noisy eigen-features. In the left and upper plots, we examine task vectors between verb past tenses and their participle (e.g., $\mathbf{went}-\mathbf{going}$). In Appendix B we show that these observations hold for other semantic binaries. (Left.) The spectrum of the Gram matrix (histogram) is well-described by a Marchenko-Pastur distribution (orange) plus an outlier “spike,” across model sizes $d$. See Section A.6 for details. (Top.) The spike corresponds to the average task vector, which comprises a few dominant eigen-features. Many of these features correspond to concepts related to history or temporal change, consistent with this semantic category. (Bottom.) We measure the strength of the spike across model size $d$ for various semantic categories. We find that the spike strength correlates strongly with the model’s ability to use the task vectors for analogy completion.
5 Discussion
@@
The primary limitation of our theory is its restriction to 1. However, we showed that our result still matches very closely with word2vec, so it is not a serious shortcoming. Empirically, we did not train and evaluate at scales that are considered state-of-the-art. We also did not compare against other word embeddings models such as GloVe (Pennington et al., 2014).
A theory of feature learning.
Our result describing an implicit bias towards low rank directly contrasts the well-studied neural tangent kernel training regime (Jacot et al., 2018; Chizat et al., 2019; Karkada, 2024). In this regime, training dynamics and generalization performance can be characterized exactly, but the models do not learn task-relevant features (Lee et al., 2019; Bordelon et al., 2020; Simon et al., 2023a). This work gives one of the first complete closed-form theories of feature learning in a minimal yet relevant natural language task. Thus, our work is a step forward in the broader project of obtaining realistic analytical solutions describing the performance of various machine learning algorithms.
Acknowledgements. DK thanks Google DeepMind for funding.
Author contributions. DK developed the analytical results, ran all experiments, and wrote the manuscript with input from all authors. JS proposed the initial line of investigation and provided insight at key points in the analysis. YB and MD helped shape research objectives and gave feedback and oversight throughout the project’s execution.
@@ A.2 Optimization
We optimize using stochastic gradient descent with no momentum nor weight decay. Each minibatch consists of 100,000 word pairs (50,000 positive pairs and 50,000 negative pairs). For benchmark evaluations (Figure 3) we use a stepwise learning rate schedule in which the base learning rate is decreased by 90% at $t=0.75t_{\mathrm{max}}$. We found that this was very beneficial, especially for QWEMs, which appear more sensitive to large learning rates. The finite-stepsize gradient descent dynamics of matrix factorization problems remains an interesting area for future research.
@@ A.3 Reweighting hyperparameters
$P_{acc}(i)=\min\left(1,\frac{10^{-5}}{P_{i}}+\sqrt{\frac{10^{-5}}{P_{i}}}\right).$ (15)
@@
Finally, Mikolov et al. (2013) draw negative samples from a different distribution. Empirically, they find that $P_{j}^{3/4}$ is particularly performant. This is equivalent to a further subsampling, exclusive to negative samples, and can be accomplished by setting
$\Psi^{-}_{ij}=\frac{P_{j}^{-1/4}}{V^{-1}\sum_{j^{\prime}}P_{j^{\prime}}^{-1/4}}.$ (17)
Note that this is an asymmetric choice, so it is not covered by 1. However, both this and the previous subsampling technique are methods for modifying and flattening the word distributions, as seen by the training algorithm. Our 1 accomplishes the same thing. We speculate that for language data, these manipulations are akin to whitening for image data.
Taken together, these give a prescription for cleanly and concisely capturing several of the implementation details of word2vec in a set of training hyperparameters. In our experiments, when training word2vec, we use all of these settings – they can be combined by multiplying the different $\Psi^{+}$ and likewise for $\Psi^{-}$. For QWEMs, we use $\Psi^{+}_{ij}=\Psi^{-}_{ij}=(P_{ij}+P_{i}P_{j})^{-1}$, and modify $\Psi^{+}$ by combining with the setting for dynamic context windows.
@@ A.4 Benchmarks
$\mathrm{acc}({\bm{W}})$ $\coloneqq\frac{1}{|\mathcal{D}|}\sum_{({\bm{a}},{\bm{b}},{\bm{a}}^{\prime},{\bm{b}}^{\prime})\in\mathcal{D}}\mathbf{1}_{\{{\bm{b}}^{\prime}\}}\left(\operatorname*{\arg\,\max}_{{\bm{w}}\in\{{\bm{W}}\}\setminus\{{\bm{a}},{\bm{b}},{\bm{a}}^{\prime}\}}{\hat{\bm{w}}}^{\top}\left(\hat{\bm{a}}^{\prime}+(\hat{\bm{b}}-\hat{\bm{a}})\right)\right)$ (18)
@@
where the 4-tuple of embeddings $({\bm{a}},{\bm{b}},{\bm{a}}^{\prime},{\bm{b}}^{\prime})$ constitute an analogy from the benchmark data $\mathcal{D}$, hats (e.g., $\hat{\bm{w}}$) denote normalized unit vectors, $\mathbf{1}$ is the indicator function, and $\{{\bm{W}}\}$ is the set containing the word embeddings. The first expression measures cosine alignment between embeddings and the “expected” representation obtained by summing the query word and the task vector. The second expression measures the degree to which four embeddings form a closed parallelogram (since they are unit vectors). The two forms of the accuracy are equivalent since
@@
We found that the accuracy metric using this selection criterion yields performance measurements that are nearly identical to those given by the standard fully-normalized accuracy metric. We conclude that the primary role of embedding normalization is to prevent longer ${\bm{w}}$’s from “winning” the $\arg\max$ just by virtue of their length. The lengths of ${\bm{a}},{\bm{b}},{\bm{a}}^{\prime}$ are only important if there is significant angular discrepancy between $(\hat{\bm{a}}^{\prime}+\hat{\bm{b}}-\hat{\bm{a}})$ and $({\bm{a}}^{\prime}+{\bm{b}}-{\bm{a}})$; in the high-dimensional regime with relatively small variations in embedding length, we expect such discrepancies to be negligible.
@@ A.5 Computational resources
Our implementations are relatively lightweight. The models can be trained on a home desktop computer with an i7 4-core processor, 32GB RAM, and a consumer-grade NVIDIA GTX 1660 graphics card, in about 2 hours. For the experiments in Figure 3, we train for 12 hours with a much lower learning rate. Our code is publicly available at https://github.com/dkarkada/qwem.
@@ A.6 Empirics of task vectors
We construct the theoretical embeddings ${\bm{W}}$ by constructing ${\bm{M}}^{\star}$ according to Equation 5, diagonalizing it, and applying 1. Due to right orthogonal symmetry, we are free to apply any right orthogonal transformation; we choose the identity as the right singular matrix, so that the components of each embedding in the canonical basis are simply its projections on the eigen-features. With this setup, it is easy to extract the embeddings resulting from a smaller model: simply truncate the extraneous columns of ${\bm{W}}$.
@@
We emphasize that, due to our Figure 2, any result comparing the final embeddings of many models of different sizes $d$ is fully equivalent to a result considering the time course of learning for a single (sufficiently large) model. Therefore, although we vary the model dimension in our plots, these can be viewed as results concerning the dynamics of learning in word embedding models.
@@
To fit the MP law to the Gram matrix of task vectors, we first subtract off the mean task vector. Then we compute the population variance of the elements of the centered matrix – we use this to set the $\sigma^{2}$ parameter. We manually fit the aspect ratio $\gamma$, and we find that the best-fitting $\hat{\gamma}=N/d_{\mathrm{eff}}$ corresponds to an effective dimension which is often much lower than the true embedding dimension $d$. We hypothesize that $d_{\mathrm{eff}}<d$ due to anisotropic noise in the task vectors. We report the $d_{\mathrm{eff}}$ resulting from the fit in Figure 4.

### v3

@@ 4 Linear Structure in Latent Space
Figure˜2 reveals that, for QWEMs, the “natural” basis of the learning dynamics is simply the eigenbasis of $\bm{M}$$^{\star}$. Figure˜3 suggests that this basis already encodes concepts interpretable to humans. These are the fundamental features learned by the model: each word embedding is naturally expressed as a linear combination of these orthogonal latents. Both early stopping and small embedding dimension regularize the model by constraining the number (but not the character) of available latents.
We may conclude that natural language contains linear semantic structure in its co-occurrence statistics, and that it is easily extracted by word embedding algorithms. It is not unreasonable to expect, then, that other semantic concepts may be encoded as linear subspaces. This idea is the linear representation hypothesis, and it motivates modern research in more sophisticated language models, including representation learning (Park et al., 2023; Jiang et al., 2024; Wang et al., 2024), mechanistic interpretability (Li et al., 2023; Nanda et al., 2023; Lee et al., 2024), and LLM alignment (Lauscher et al., 2020; Zou et al., 2023; Li et al., 2024). To make these efforts more precise, it is important to develop a quantitative understanding of these linear representations in simple models. Our ˜1 provides a new lens for this analysis: we may gain insight by studying the properties of $\bm{M}$$^{\star}$.
In word embedding models, a natural category of abstract linear representations are the difference vectors between semantic pairs, e.g., $\{{\bm{r}}^{(n)}\}_{n}=\{\mathbf{man}-\mathbf{woman},\mathbf{king}-\mathbf{queen},\dots\}$ for gender binaries. If the ${\bm{r}}^{(n)}$ are all approximately equal, then the model can effectively complete analogies via vector addition (see Section˜A.4). Following Ilharco et al. (2022), we call the ${\bm{r}}^{(n)}$ task vectors.
@@
We emphasize that, due to our Figure˜2, any result comparing the final embeddings of many models of different sizes $d$ is fully equivalent to a result considering the time course of learning for a single (sufficiently large) model. Therefore, although we vary the model dimension in our plots, these can be viewed as results concerning the dynamics of learning in word embedding models.
@@
Task vectors derived from the analogy dataset are neither strongly aligned with a single model eigen-feature, nor are they random vectors. Instead, they exhibit localization: a handful of the top eigen-features are primarily responsible for the task vector. In some cases, these dominant eigen-features are interpretable and clearly correlate with the abstract semantic concept associated with the task vector (Figure˜4).
@@
To study the geometry of the task vectors within a class of semantic binaries, we consider stacking task vectors to produce the matrix ${\bm{R}}_{d}\in\mathbb{R}^{N\times d}$, where $N$ is the number of word pairs and $d$ is the embedding dimension. We note that ${\bm{R}}_{d}$ can be computed in closed form using our ˜1. If all the task vectors align (as desired for analogy completion), then ${\bm{R}}_{d}$ will be a rank-1 matrix; if the task vectors are all unrelated, then ${\bm{R}}_{d}$ will have a broad spectrum. This observation suggests that a useful object to study is the empirical spectrum of the Gram matrix ${\bm{G}}_{d}\coloneqq{\bm{R}}_{d}{\bm{R}}_{d}^{\top}\in\mathbb{R}^{N\times N}$.
We find that this spectrum is very well-described by the spiked covariance model, a well-known distribution of random matrices. In the model, one studies the spectrum of ${\bm{Z}}=\bm{\Xi}{\bm{\Xi}}^{\top}+\mu{\bm{a}}{{\bm{a}}}^{\top}$, where $\bm{\Xi}$ is a random matrix with i.i.d. mean-zero entries with variance $\sigma^{2}$ and $\mu{\bm{a}}{{\bm{a}}}^{\top}$ is a deterministic rank-1 perturbation with strength $\mu$. If $\mu$ is sufficiently large compared to $\sigma$, then in the asymptotic regime the spectrum of $\bm{Z}$} is known to approach the Marchenko-Pastur distribution with a single outlier eigenvalue (Baik et al., 2005). We consistently observe this structure in the real empirical data, across both model dimension and semantic families (Figure˜4). We note that $N\approx 30$ is fairly small, and so it is somewhat surprising that results from the asymptotic regime visually appear to hold.
@@
Figure 4: Models build linear representations from a few informative and many noisy eigen-features. In the left and upper plots, we examine task vectors between verb past tenses and their participle (e.g., $\mathbf{went}-\mathbf{going}$). In Appendix˜B we show that these observations hold for other semantic binaries. (Left.) The spectrum of the Gram matrix (histogram) is well-described by a Marchenko-Pastur distribution (orange) plus an outlier “spike,” across model sizes $d$. See Section˜A.6 for details. (Top.) The spike corresponds to the average task vector, which comprises a few dominant eigen-features. Many of these features correspond to concepts related to history or temporal change, consistent with this semantic category. (Bottom.) We measure the strength of the spike across model size $d$ for various semantic categories. We find that the spike strength correlates strongly with the model’s ability to use the task vectors for analogy completion.
5 FAQ
Can the assumptions be relaxed?
Our theoretical results require only four assumptions: quartic approximation of the loss, our ˜1, small initial weights, and population gradient flow. The latter two are technical conveniences that simplify the analysis; Figure˜2 suggests that they may be relaxed (or possibly eliminated) with additional effort. The first two are genuine approximations of the word2vec algorithm; their validity is supported by our empirical checks (see Figures˜1, 3, A.3 and 5). To further understand why these two approximations are technically useful, and what happens if they are relaxed, we perform empirical ablation tests in Figure˜7.555We note that these auxiliary experiments use a different training corpus, model size, and hyperparameter choices, indicating that our theory is not sensitive to these choices.
The clear stepwise learning and decoupled dynamics in our theory originate from ˜1. This condition turns the weighted factorization problem into an unweighted factorization. As a consequence, the singular value/vector dynamics quickly become “untangled” in the early stage of training; see ˜1 and the subsequent discussion. Somewhat surprisingly, this simplification is “close enough” to word2vec. Though the word2vec singular vectors do exhibit mild “mixing” in time (Figure˜2), the overall learning dynamics are well-described by the simplified setting. This agreement is partly because the recommended word2vec hyperparameters approximate satisfy ˜1 (see Section˜A.3). Our results suggest that future efforts to understand more complex learning systems may find purchase in identifying useful approximations of this kind.
Why do QWEM factorizations describe word2vec better than PMI factorizations?
The crux of the issue lies in adequately handling the rank constraint. The idea behind factorizing PMI via SVD is to first solve the unconstrained minimizer of the loss (i.e., the PMI matrix) and then, at the end, apply the rank constraint by choosing the closest (in Frobenius norm) rank-$d$ matrix. This does not work well because the loss basin is extremely wide and shallow, so the global unconstrained minimum is actually very far from where the model actually ends up via gradient descent in finite time. We take a different approach: by first approximating the loss landscape itself, we can account for the rank constraint throughout the entire optimization trajectory. As a result, our prediction for what word2vec learns is significantly more accurate.
Why does decreasing SNR sometimes still yield high performance in Figure˜4?
This is an ancillary effect of using top-1 accuracy as the performance metric. As a concrete example, consider the analogy “France : Paris :: Japan : Tokyo,” satisfying $\texttt{japan}+(\texttt{paris}-\texttt{france})\approx\texttt{tokyo}$. As the effective embedding dimension increases, the SNR may decrease; at the same time, the typical separation between embeddings increases (due to the increased available volume in latent space). This means that there are fewer nearby “competitors” for tokyo. In the large $d$ regime, we find that the embeddings are sufficiently spaced, and tokyo is still the nearest embedding to $\texttt{japan}+(\texttt{paris}-\texttt{france})$ despite an increase in absolute error (see Figure˜8). Thus, top-1 performance does not degrade with decreasing SNR. We note that unintuitive effects associated with using top-$k$ accuracy have been observed in LLMs as well (Schaeffer et al., 2024).
How might one apply these results to other self-supervised learning tasks?
Since our results are distribution agnostic, ˜1 can be extended to establish an explicit equivalence between self-supervised contrastive learning and supervised matrix factorization. In particular, if the contrastive objective has the functional form $\mathcal{L}({\bm{W}})=\mathbb{E}_{\text{positive pairs}}[f^{+}({\bm{w}}^{T}{\bm{w}}^{\prime})]+\mathbb{E}_{\text{negative pairs}}[f^{-}({\bm{w}}^{T}{\bm{w}}^{\prime})]$ for some differentiable $f^{+}(\cdot)$ and $f^{-}(\cdot)$, then one can simply take the quadratic Taylor approximation, complete the square, and obtain a target matrix in terms of the positive and negative distributions of the learning problem. See Section˜D.2 for a demonstration of this idea for the SimCLR loss. To obtain closed-form learning dynamics, one needs to find an equivalent of our ˜1. This may be challenging depending on the form of the input data; for instance, when the inputs are not one-hot, it may require a whitened data covariance.
What do these results tell us about feature learning?
Two operational notions of feature learning are 1) optimization trajectories must escape the local vicinity of the (typically small) initialization (Yang & Hu, 2021; Jacot et al., 2021; Zhu et al., 2022; Atanasov et al., 2024; Kunin et al., 2025), and 2) learned weight matrices must project data onto target-relevant subspaces (Damian et al., 2022; Radhakrishnan et al., 2024). Our message complements this line of research by offering a practical yet solvable setting in which both behaviors appear: word2vec escapes its near-isotropic initialization region to learn a dense compression of the relative excess co-occurrence probability $\bm{M}$$^{\star}$, thereby aligning the embedding geometry with the most salient semantic linear subspaces. Our result is thus a step forward in the broader scientific project of obtaining quantitative, predictive descriptions of learning in practical algorithms.
@@ 5 FAQ
Our results are limited to the symmetric setup (tied encoder/decoder weights, i.e., ${\bm{W}}={\bm{W}}^{\prime}$). We did not train and evaluate at scales that are considered state-of-the-art, nor did we compare against other embedding models such as GloVe (Pennington et al., 2014).
Author contributions. DK developed the analytical results, ran all experiments, and wrote the manuscript with input from all authors. JS proposed the initial line of investigation and provided insight at key points in the analysis. YB and MD helped shape research objectives and gave oversight and feedback throughout the project’s execution.
Acknowledgements. DK is grateful to Daniel Kunin and Carl Allen for early conversations about learning linear analogy structure from small initialization. DK thanks Google DeepMind and BAIR for funding support and compute access. MRD thanks the U.S. Army Research Laboratory and the U.S. Army Research Office for supporting this work under Contract No. W911NF-20-1-0151.
@@ A.2 Optimization
We optimize using stochastic gradient descent with no momentum nor weight decay. Each minibatch consists of 100,000 word pairs (50,000 positive pairs and 50,000 negative pairs). For benchmark evaluations (Figure˜3) we use a stepwise learning rate schedule in which the base learning rate is decreased by 90% at $t=0.75t_{\mathrm{max}}$. We found that this was very beneficial, especially for QWEMs, which appear more sensitive to large learning rates. The finite-stepsize gradient descent dynamics of matrix factorization problems remains an interesting area for future research.
@@ A.3 Reweighting hyperparameters
$P_{acc}(i)=\min\left(1,\frac{10^{-3}}{P_{i}}+\sqrt{\frac{10^{-3}}{P_{i}}}\right).$ (15)
@@
Finally, Mikolov et al. (2013) draw negative samples from a different distribution (empirically, they find that $P_{j}^{3/4}$ is particularly performant) and upweight the negative sampling term by a hyperparameter $k$ (they recommend $k\approx 5$ for large corpora). This is equivalent to setting
$\Psi^{-}_{ij}=\frac{kP_{j}^{-1/4}}{V^{-1}\sum_{j^{\prime}}P_{j^{\prime}}^{-1/4}}.$ (17)
Note that this is an asymmetric choice, so it is not covered by ˜1. However, both this and the previous subsampling technique are methods for modifying and flattening the word distributions, as seen by the training algorithm. Our ˜1 accomplishes the same thing. We speculate that these manipulations accomplish for language data what spectral whitening does for image data.
Taken together, these give a prescription for cleanly and concisely capturing several of the implementation details of word2vec in a set of training hyperparameters. In our experiments, when training word2vec, we use all of these settings – they can be combined by multiplying the different $\Psi^{+}$ and likewise for $\Psi^{-}$. For QWEMs, we use $\Psi^{+}_{ij}=\Psi^{-}_{ij}=(P_{ij}+P_{i}P_{j})^{-1}$, and modify $\Psi^{+}$ by combining with the setting for dynamic context windows. For fair comparison, we use $k=1$ for the SGNS experiments in the main text.
@@ A.4 Benchmarks
[…]
@@
where the 4-tuple of embeddings $({\bm{a}},{\bm{b}},{\bm{a}}^{\prime},{\bm{b}}^{\prime})$ constitute an analogy from the benchmark data $\mathcal{D}$, hats (e.g., $\hat{\bm{w}}$) denote normalized unit vectors, $\mathbf{1}$ is the indicator function, and $\{{\bm{W}}\}$ is the set containing the word embeddings. The first expression measures cosine alignment between embeddings and the “expected” representation obtained by summing the query word and the task vector. The second expression measures the degree to which four embeddings form a closed parallelogram (Rumelhart & Abrahamson, 1973). The two forms of the accuracy are equivalent since
@@
[…]
@@ A.5 Computational resources
Our implementations are relatively lightweight. The models can be trained on a home desktop computer with an i7 4-core processor, 32GB RAM, and a consumer-grade NVIDIA GTX 1660 graphics card, in about 2 hours. For the experiments in Figure˜3, we train for 12 hours with a much lower learning rate. Our code is publicly available at https://github.com/dkarkada/qwem.
@@ A.6 Empirics of task vectors
We construct the theoretical embeddings $\bm{W}$} by constructing $\bm{M}$$^{\star}$ according to Equation˜5, diagonalizing it, and applying ˜1. Due to right orthogonal symmetry, we are free to apply any right orthogonal transformation; we choose the identity as the right singular matrix, so that the components of each embedding in the canonical basis are simply its projections on the eigen-features. With this setup, it is easy to extract the embeddings resulting from a smaller model: simply truncate the extraneous columns of $\bm{W}$}.
@@
To fit the MP law to the Gram matrix of task vectors, we first subtract off the mean task vector. Then we compute the population variance of the elements of the centered matrix – we use this to set the $\sigma^{2}$ parameter. We manually fit the aspect ratio $\gamma$, and we find that the best-fitting $\hat{\gamma}=N/d_{\mathrm{eff}}$ corresponds to an effective dimension which is often much lower than the true embedding dimension $d$. We hypothesize that $d_{\mathrm{eff}}<d$ due to anisotropic noise in the task vectors. We report the $d_{\mathrm{eff}}$ resulting from the fit in Figure˜4.